OpenStudy (anonymous):

It lies above but I say none of them holds true: Which test point holds true for the inequality 32y−2x≥1? Where does the shaded area lie for this inequality? The test point (1/4, 3/4) (1/4, 1) (1, 0) (1, 1/4) (4, 1) holds true for this inequality. The shaded area for the inequality lies above or below the boundary line.

jimthompson5910 (jim_thompson5910):

Plug each (x,y) point into the inequality and see which point makes the inequality true.

jimthompson5910 (jim_thompson5910):

For example, testing (x,y) = (1,0) gives 32y - 2x >= 1 32(0) - 2(1) >= 1 0 - 2 >= 1 -2 >= 1... false Since -2 >= 1 is false, this means 32y - 2x >= 1 is false when (x,y) = (1,0)

OpenStudy (acxbox22):

@perl

OpenStudy (anonymous):

It's (1/4,3/4)

jimthompson5910 (jim_thompson5910):

what else works?

OpenStudy (anonymous):

(1/4,1)

jimthompson5910 (jim_thompson5910):

hmm I'm thinking there might be a typo

OpenStudy (anonymous):

(1/4,3/4)(1/4,1)(1,0)(1,1/4)(4,1)

jimthompson5910 (jim_thompson5910):

only one point doesn't work while the others do

jimthompson5910 (jim_thompson5910):

This is what I get with geogebra